Let A be a 2 x 2 matrix. Show that some non-trivial linear combination of A^4, A^3, A^2, A. and I2 is equal 0. Generalize to n x n matrices. Note that I2 is 2 x 2 identity matrix.


1
Expert's answer
2021-09-30T00:30:14-0400

Let "A=I=\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}" "(2\times 2 \ matrix)"

"\therefore A^4=A^3=A^2=I=\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

Linear Combination:

"aA^4+bA^3+cA^2+dA+eI^2=0"

"\Rightarrow (a+b+c+d+e)\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}=0"

"\therefore a+b+c+d+e=0"


Now, solving for "(n\times n \ matrix):"


Let "A= I_n=\begin{bmatrix}
 1 & 0 & 0& 0& 0& 0& 0 .......& 0\\
 0 & 1 & 0& 0& 0& 0& 0.......& 0\\
0 & 0 & 1& 0& 0& 0& 0.......& 0\\
.
\\
.
\\
.
\\
.
0 & 0 & 0& 0& 0& 0& 0.......& 1\\


\end{bmatrix}"

"\therefore A^4=A^3=A^2=A= I_n=\begin{bmatrix}
 1 & 0 & 0& 0& 0& 0& 0 .......& 0\\
 0 & 1 & 0& 0& 0& 0& 0.......& 0\\
0 & 0 & 1& 0& 0& 0& 0.......& 0\\
.
\\
.
\\
.
\\
.
0 & 0 & 0& 0& 0& 0& 0.......& 1\\


\end{bmatrix}"

Linear Combination:

"aA^4+bA^3+cA^2+dA+eI_n^2=0"

"\Rightarrow (a+b+c+d+e)\begin{bmatrix}
 1 & 0 & 0& 0& 0& 0& 0 .......& 0\\
 0 & 1 & 0& 0& 0& 0& 0.......& 0\\
0 & 0 & 1& 0& 0& 0& 0.......& 0\\
.
\\
.
\\
.
\\
.
0 & 0 & 0& 0& 0& 0& 0.......& 1\\


\end{bmatrix}=0"

"\therefore a+b+c+d+e=0"